A train travels 360km at a uniform speed. If the speed had been 5km/h more, it would have taken 1 hour less for the same journey. Find the speed of the train
Answer by mananth(16946)

distance = 360 km
..
let speed be x
then time taken = d/r
t=360/x
..
if speed is increased by 5 km/h
speed will be (x+5)km/h
..
distance is the same
t= 360/(x+5)
..
time with original speed - time with increased speed = 1
..
360/x - 360/(x+5)=1
LCD= x(x+5)
(360(x+5)-360x) /x(x+5)=1
360x+1800-360x=x(x+5)
x^2+5x = 1800
x^2+5x-1800=0
x^2+45x-40x-1800=0
x(x+45)-40(x+45)=0
(x-40)(x+45)=0
x= 40 mph the original speed.
...
m.ananth@hotmail.ca